# Learning graph geometry and topology using dynamical systems based message-passing

### v3

@@ 3.1 Motivation: Connection between dynamics and topological properties
Due to Proposition 3.1 we make the assumption that $G$ is connected in the following sections. However, we note that many of the results still apply to disconnected graphs with suitable modifications.
We now discuss several desirable attributes of the heat, wave, and Sprott dynamics which stem from the structure of the differential equation and their corresponding solutions. The dynamics we consider are canonical and capture a wide range of solution behavior, thereby enabling the results of this work to set the groundwork for future dynamics based message passing schemes.
@@ 3.1.1 Heat Dynamics
Finally, we restate a known results that provide additional foundation linking the behavior of the heat equation solutions to graph topology. Lemma 1 of Crane et al. [12] shows that the heat equation encodes shortest path distances $d(u,v)$ between nodes $u,v$ on the graph:
@@
We additionally note that the ability of the heat equation to capture intrinsic graph distances links it to graph Ollivier-Ricci curvature $\kappa$ and provide a detailed statement in the appendix.
@@ 3.1.2 Wave Dynamics
Eqn. 8, shows that unlike the heat energy which is non-increasing, the wave energy will oscillate over time. Thus, the wave equation may be well suited for capturing global geometric features which require long-range interactions. The form that the solution to the wave equation takes on makes it particularly well suited to capturing aspects of graph topology. In particular, we find that it provides global information from local observation.
@@
The graph spectra encodes a wide range of graph invariants and properties. The above proposition demonstrates that the solutions to the wave equation relate to graph spectral properties, and that this entire information is contained in the solutions of the wave equation at each node. As an example of how this may be related to graph topology, for instance in a cycle graph, the eigenvalues determine the length of the cycle. More generally, when the graph is not a cycle but contains a cyclic subgraphs as a prominent topological feature, this proposition provides some intuition for why the wave-equation is well suited to pick up that a node belongs to a cycle and recover dimension 1 homology.
3.1.3 Chaotic dynamics
Based on choice of $b=0.25$, it was observed in Sprott [37] that certain networks exhibited weakly chaotic (small positive Lyapunov exponents) dynamics. Weakly dynamic systems retain some memory but can explore much larger portions of the state space, thereby creating highly complex representations. In Lai & Winslow [24] a spatio-temporal chaotic dynamical system demonstrated highly sensitive dependence to initial condition in phase space and in the parameter space of this system. Thus chaotic network dynamics may be useful in situations where it is important to distinguish small differences in graph structure and data.
@@ 3.2 The DYMAG Architecture
Depending on if the task is defined at the node or graph level, we pass a flattened version of the node level representation (Eqn. 9) or the graph level representation (Eqn. 10) to a multi-layer perceptron (MLP). This is the trainable part of DYMAG. The construction of DYMAG is flexible and amenable to incorporation with other models. We leave such exploration to future work. Our code is publicly available [^1] https://anonymous.4open.science/r/DYMAG-196E/, and additional implementation details are described in Section B of the appendix.
@@ Appendix A Proofs of Theoretical Results
The first inequality in Eqn. 13 follows by noting:
@@
Eqn. 14 immediately follows Eqn. 15 and the fact that $0=\lambda_{1}<\lambda_{2}\leq\ldots\leq\lambda_{n}$. ∎
@@
The proof for Eqn. 8 is similar to the proof of Eqn. 12.
@@
Now, let $\epsilon>0$ be a degree of percision and choose $K$ such that $\frac{1}{K}<\epsilon$. Approximate
@@
A.7 The proof of Proposition A.2
@@
Substituting $\lambda=1-\mu$ in Eqn. 18 links the eigenvalues of $U^{t}_{H}$ and $P$ by
@@
A.8 Relation of the heat equation to Ollivier-Ricci Curvature
The Ollivier-Ricci curvature is a discrete notion of curvature, meant to paralell the tradition notion of Ricci curvature in Riemmanian geometry. It is defined by
$\kappa(v,w)=1-W_{1}(\mu_{v},\mu_{w})/d(v,w),$
The following result from Münch & Wojciechowski [27] related $\kappa$ to the solution of the heat equation.
Proposition A.4 (Relation to Ollivier-Ricci curvature, (Theorem 5.8 of [27])).

### v4

@@ 3.1 Motivation: Connection between dynamics and topological properties
Hence methods building upon dynamics become endowed with the ability to discern the connected components of a graph. For simplicity of notation and the establishment of proposition 3.1 we make the assumption that $G$ is connected in the following sections. However, we note that many of the results still apply to disconnected graphs with suitable modifications.
We now discuss several desirable attributes of the dynamics which stem from the structure of the differential equation and their corresponding solutions. The dynamics we consider are canonical and capture a wide range of solution behavior, thereby enabling the results of this work to set the groundwork for future dynamics based message passing schemes.
@@ 3.1.1 Heat Dynamics
Finally, we restate two known results that provide additional foundation linking the behavior of the heat equation solutions to graph topology. Lemma 1 of Crane et al. [12] shows that the heat equation encodes shortest path distances $d(u,v)$ between nodes $u,v$ on the graph:
@@
We additionally note that the ability of the heat equation to capture intrinsic graph distances links it to graph Ollivier-Ricci curvature $\kappa$. The Ollivier-Ricci curvature is a discrete notion of curvature, meant to paralell the tradition notion of Ricci curvature in Riemmanian geometry. It is defined by
$\kappa(v,w)=1-W_{1}(\mu_{v},\mu_{w})/d(v,w),$
The following result from Münch & Wojciechowski [26] relates $\kappa$ to the solution of the heat equation.
Proposition 3.6 (Relation to Ollivier-Ricci curvature, (Theorem 5.8 of [26])).
Hence heat dynamics are closely related to various geometric properties intrinsic to the graph, including notions of distance and curvature.
@@ 3.1.2 Wave Dynamics
Eqn. 9, shows that unlike the heat energy which is non-increasing, the wave energy will oscillate over time. Thus, the wave equation may be well suited for capturing global geometric features which require long-range interactions. The form that the solution to the wave equation takes on makes it particularly well suited to capturing aspects of graph topology. In particular, we find that it provides global information from local observation.
@@
The graph spectra encodes a wide range of graph invariants and properties. The above proposition demonstrates that the solutions to the wave equation relate to graph spectral properties, and that this entire information is contained in the solutions of the wave equation at each node. The following corollary serves as an example of how this may be related to graph topology.
Corollary 3.9 (Cycle length).
The size of a cycle graph can be determined from finitely many samples of the wave equation at a single node $v$.
More generally, when the graph is not a cycle but contains a cyclic subgraphs as a prominent topological feature, this proposition provides some intuition for why the wave-equation is well suited to pick up that a node belongs to a cycle and recover dimension 1 homology.
@@ 3.2 The DYMAG Architecture
Depending on if the task is defined at the node or graph level, we pass a flattened version of the node level representation (Eqn. 10) or the graph level representation (Eqn. 11) to a multi-layer perceptron (MLP). This is the trainable part of DYMAG. The construction of DYMAG is flexible and amenable to incorporation with other models. We leave such exploration to future work. Our code is publicly available [^1] https://anonymous.4open.science/r/DYMAG-196E/, and additional implementation details are described in Section B of the appendix.
@@ Appendix A Proofs of Theoretical Results
The first inequality in Eqn. 14 follows by noting:
@@
Eqn. 15 immediately follows Eqn. 16 and the fact that $0=\lambda_{1}<\lambda_{2}\leq\ldots\leq\lambda_{n}$. ∎
@@
The proof for Eqn. 9 is similar to the proof of Eqn. 13.
@@
Now, let $\epsilon>0$ be a degree of precision and choose $K$ such that $\frac{1}{K}<\epsilon$. Approximate
@@
A.7 The proof of Corollary 3.9
Proof.
Denote $C_{n}$ the cycle graph with vertices numbered $10,\ldots,n-$ 1 and edges $(v,v+1)$ modulo $n$. Since $C_{n}$ is 2-regular, the unnormalized and normalized Laplacian differ only by a constant multiple of 2. Therefore, without loss of generality we will focus on the unnormalized Laplacian.
It is known that an orthogonal eigenbasis is given by $\{\phi_{k}\}_{k=0}^{k=\lfloor n/2\rfloor}\cup\{\psi_{k}\}_{k=1}^{\lceil(n-1)/2\rceil}$ defined:
$\phi_{k}(v)=\cos\left(\frac{2\pi kv}{n}\right),\qquad\psi_{k}(v)=\sin\left(\frac{2\pi kv}{n}\right),$
where the corresponding eigenvalues are given by
$\lambda_{k}=2-2\cos\left(\frac{2\pi k}{n}\right)=4\sin^{2}\left(\frac{\pi k}{n}\right)$ (18)
Each of the $\lambda_{k}$ are uniquely determined by our measurements by Lemma 3.8, and
$n=\frac{2\pi}{\cos^{-1}\left(\frac{2-\lambda_{1}}{2}\right)}.$
Thus $n$ is uniquely determined by our measurements. ∎
A.8 The proof of Proposition A.2
@@
Substituting $\lambda=1-\mu$ in Eqn. 20 links the eigenvalues of $U^{t}_{H}$ and $P$ by